11.1A survey asked, How many tattoos do you currently have on your body? Of the 1211 males surveyed, 193 responded that they haA physical therapist wants to determine the difference in the proportion of men and women who participate in regular sustaineIn randomized, double-blind clinical trials of a new vaccine, children were randomly divided into two groups. Subjects in gro
Answer #1

Thursday

30/07/2020

Standard errors = Vp1(1-p1/n1+p2(1-p2/n2 = V0.0001+0.0001 = 0.0148 x2 = 147 n2 = 1089 p2 = x2/n2 = 147 / 1089 Given that x1 =

pls upvote...!!!